OpenStudy (anonymous):

S varies directly as T. If S is 20 when T is 4, then T is____ a0 when S is 30.

OpenStudy (anonymous):

HELP ME PLEASE!!!!!!!!!

OpenStudy (anonymous):

This is a ratio problem. S is 20 when T is 4 so when S moves to 30 then T moves to x. Looks like this; 20 / 4 = 30 / x. Now you cross multiply 20x = 120. Divide by 20 on each side to get: x = 6.

OpenStudy (anonymous):

tnx :)

OpenStudy (anonymous):

Ur welcom (: